Alex, a rescue pilot drops a survival kit while her plane is flying horizontally at an altitude of 1800.0 m with a forward velocity of 140.0 m/s. A group of lost Antarctica explorers is below on the ground.
How far in front of the explorers should Alex drop the supplies?
1
Expert's answer
2019-10-03T09:28:27-0400

The time of flight:


"t=\sqrt{\frac{2h}{g}}"

"d=vt=v\sqrt{\frac{2h}{g}}=140\sqrt{\frac{2(1800)}{9.8}}=2683\ m"